OpenStudy (anonymous):

@order @badreferences - (2x)^3 = what?

OpenStudy (anonymous):

\(\left(ab\right)^n=a^nb^n\) Can't believe I typed this out twice today.

OpenStudy (dumbcow):

note: look 2 post down

OpenStudy (anonymous):

Note: WolframAlpha http://www.wolframalpha.com/ is a lovely thing.

OpenStudy (anonymous):

Just don't get addicted.

OpenStudy (anonymous):

distribute the exponent to the terms inside the parenthesis. .

OpenStudy (dumbcow):

wolfram is great

OpenStudy (anonymous):

is the answer, 8x or 8x^3 ?

OpenStudy (anonymous):

The latter; the operation of raising something to a power distributes multiplicatively.